It takes Meihua two hours more to complete a 50-km journey than it takes Ailin to complete a 40-km journey.If the average speed of Meihua for the journey is 5 km/h less than ailin, calculate the average speed for each girl.
Answer by mananth(16946)

It takes Meihua two hours more to complete a 50-km journey than it takes Ailin to complete a 40-km journey.If the average speed of Meihua for the journey is 5 km/h less than ailin, calculate the average speed for each girl.
Let ailin's speed be x
mehua's speed will be x-5
Time taken by ailin to cover 40 km= 40/x
.
Time taken by mehua to cover 50 km = 50/x-5
Mehua takes 2 hours more
.
50/x-5 - 2 = 40/x
50/x-5 -40/x = 2
50x-40(x-5) / x(x-5)=2
50x-40x+200= 2x(x-5)
50x-40x+200= 2x^2 -10x
2x^2-20x-200=0
x^2-10x-100=0
x1,x2 the roots of the equation
will be x1 = 10+sqrt(100+400) / 2
x1= 16.18 mph the speed of ailin
x2 =-6.18 which is absurd.
Mehua's apeed will be 11.18 mph